# Furstenberg set problem and exceptional set estimate in prime fields: dimension two implies higher dimensions

### v1

@@ 4. Furstenberg set problem: Lower bound I
Suppose $E=\bigcup_{V\in\mathcal{V}}Y(V)$. Note that each $x\in E$ can lie in at most #$\#G(k,\mathbb{F}_{p}^{n})\sim p^{k(n-k)}$ different $V$. We have
#$\#E\cdot p^{k(n-k)}\geq\sum_{V\in\mathcal{V}}\#Y(V)\gtrsim p^{t}.$
Hence, #$\#E\gtrsim p^{t-k(n-k)}$. Also, trivially, #$\#E\geq 1$. ∎
Proposition 28.
@@
By the definition of $(s,0;n,k)$-set, #$\#E\geq\#Y(V)\gtrsim p^{s}$. ∎
In this section, we prove Theorem 7 for $(n,k)=(k+1,k)$. The proof is by induction on $k$. The goal of this section is to prove the following result.
Proposition 29.
Let $k\in\mathbb{N}$ and $k\geq 2$. Suppose for admissible $(s,t;k,k-1)$, we know
(27) $\inf_{E:\ (s,t;k,k-1)\textup{-set}}\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;k,k-1)-\varepsilon}$
@@
(28) $\inf_{E:\ (s,t;k+1,k)\textup{-set}}\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;k+1,k)-\varepsilon},$
@@
We first prove the following lemma.
Lemma 30.
Assume (27) holds. Suppose $(s,t;k+1,k)$ is admissible. Then for any $(s,t;k+1,k)$-set $E$, there exist numbers $t_{1},t_{2},s_{1},s_{2},u,v$, such that
(29) $\begin{cases}1\leq p^{t_{1}}\leq p^{k-1},1\leq p^{t_{2}}\leq p^{2};\\ 1\leq p^{s_{1}}\leq p,1\leq p^{s_{2}}\leq p^{s};\\ p^{s_{1}}\lesssim p^{u}\leq p,1\leq p^{v}\leq p;\end{cases}$
@@
(30) $\begin{cases}p^{t_{1}+t_{2}}\gtrapprox p^{t};\\ p^{s_{1}+s_{2}}\gtrapprox p^{s};\\ p^{u+v}\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{1},t_{2};2,1)-\varepsilon};\end{cases}$
@@
(31) #$\#E\gtrsim_{\varepsilon}p^{u+\max\{\mathbf{F}(s_{2},t_{1}+v;k,k-1),s_{2}+v\}-\varepsilon}.$
@@
where $\mathcal{V}\subset A(k,\mathbb{F}_{p}^{k+1})$ with #$\#\mathcal{V}\sim p^{t}$, and $Y(V)\subset V$ with #$\#Y(V)\sim p^{s}$.
@@
Now, we partition $\mathcal{V}$ according to $G(k-1,\mathbb{F}_{p}^{k}\times\mathbf{0})$. Since each $V\in\mathcal{V}$ is not parallel to $\mathbb{F}_{p}^{k}$ and the ambient space is $\mathbb{F}_{p}^{k+1}$, $V$ must intersect $\mathbb{F}_{p}^{k}\times\mathbf{0}$ at a $(k-1)$-plane. For each $W\in G(k-1,\mathbb{F}_{p}^{k}\times\mathbf{0})$, we define $\mathcal{V}_{W}$ to be the set of $V\in\mathcal{V}$ such that $V\cap\mathbb{F}_{p}^{k}$ is parallel to $W$. We obtain a partition
@@
First pigeonholing. We assume there exists $\mathcal{W}\subset G(k-1,\mathbb{F}_{p}^{k}\times\mathbf{0})$ with #$\#\mathcal{W}\sim p^{t_{1}}$ such that for each $W\in\mathcal{W}$, #$\#\mathcal{V}_{W}\sim p^{t_{2}}$. Here, $1\leq p^{t_{1}}\leq p^{k-1},1\leq p^{t_{2}}\leq p^{2}$ are dyadic numbers and
(32) $p^{t_{1}}p^{t_{2}}\gtrapprox p^{t}.$
(Recall $A\gtrapprox B$ means $A\gtrsim(\log p)^{C}B$.) We still use $\mathcal{V}$ to denote $\bigcup_{W\in\mathcal{W}}\mathcal{V}_{W}$. And we see that
#$\#\mathcal{V}\gtrapprox p^{t}.$
Second pigeonholing. By properly choosing the coordinate and throw away $\lesssim 1$ fraction of $W$ in $\mathcal{W}$, we can assume each $W$ in $\mathcal{W}$ satisfies $W\cap(\mathbb{F}_{p}\times\mathbf{0}^{k})=\mathbf{0}^{k+1}$. Denote $\widetilde{\mathbb{F}_{p}^{2}}=\mathbb{F}_{p}\times\mathbf{0}^{k-1}\times\mathbb{F}_{p}$. If $V\in\mathcal{V}_{W}$ and $W\in\mathcal{W}$, then $V\cap\widetilde{\mathbb{F}_{p}^{2}}$ is a line. We denote it by $\ell_{V}:=V\cap\widetilde{\mathbb{F}_{p}^{2}}$. Note that $\ell_{V}$ is not parallel to the first coordinate in $\widetilde{\mathbb{F}_{p}^{2}}$.
@@
Noting that #$\#Y(V)\sim p^{s}$, by pigeonholing, we find a subset $Y(\ell_{V})\subset\ell_{V}$, so that #$\#Y(\ell_{V})\sim s_{1}(V)$, and for each $x\in Y(\ell_{V})$, #$\#Y(V)\cap(x+W)\sim s_{2}(V)$. Here, $1\leq p^{s_{1}(V)}\leq p,1\leq p^{s_{2}(V)}\leq p^{s}$ are dyadic numbers and satisfy
@@
$Y(V,x)$ is called the horizontal $x$-slice. (Note that there is no confusion, since $V$ uniquely determines $W$.) Thus, we have
(33) $Y(V)=\bigsqcup_{x\in Y(\ell_{V})}Y(V,x).$
@@
We summarize what we have done so far. We can redefine $\mathcal{W}$, $\mathcal{V}_{W}$ and $\mathcal{V}$, so that $\mathcal{V}=\bigcup_{W\in\mathcal{W}}\mathcal{V}_{W}$ and the following uniformity holds.
#$\#\mathcal{W}\gtrapprox p^{t_{1}}.$
#$\#\mathcal{V}_{W}\gtrapprox p^{t_{2}},\textup{~{}for~{}}W\in\mathcal{W}.$
@@
(34) $Y(V)=\bigsqcup_{x\in Y(\ell_{V})}Y(V,x),$
with #$\#Y(\ell_{V})\sim p^{s_{1}}$ and for each $x\in Y(\ell_{V})$, #$\#Y(V,x)\sim p^{s_{2}}$.
Fix a $W\in\mathcal{W}$. We want to estimate $\bigcup_{V\in\mathcal{V}_{W}}Y(\ell_{V})(\subset\widetilde{\mathbb{F}_{p}^{2}})$. This is some sort of Furstenberg set in two-dimensional space. Also, an important preperty is that each $\ell_{V}$ is not parallel to the first coordinate. We will exploit some “quasi-product” structure of this set.
Lemma 31.
Assume Conjecture 4 holds. Suppose $E=\bigcup_{\ell\in\mathcal{L}}Y(\ell)$ is a $(s_{1},t_{2};2,1)$-set. Also, we assume each $\ell\in\mathcal{L}$ is not parallel to $\mathbb{F}_{p}\times\mathbf{0}$. Then there exist dyadic numbers $p^{u},p^{v}$ such that
@@
for any $\varepsilon>0$. Also, $E$ contains a subset of form
@@
where $\Xi\subset\mathbf{0}\times\mathbb{F}_{p}$ with #$\#\Xi\sim p^{u}$, and $Y_{\xi}\subset E\cap\bigg{(}\xi+(\mathbb{F}_{p}\times\mathbf{0})\bigg{)}$ with #$\#Y_{\xi}\sim p^{v}$.
@@
For $\xi\in\mathbf{0}\times\mathbb{F}_{p}$, define $E(\xi):=E\cap\bigg{(}\xi+(\mathbb{F}_{p}\times\mathbf{0})\bigg{)}$. By pigeonholing on horizontal slices of $E$, we can find $\Xi\subset\mathbf{0}\times\mathbb{F}_{p}$ so that for all $\xi\in\Xi$, $E(\xi)$ have comparable cardinality. We may denote #$\#\Xi\sim p^{u},\#E(\xi)\sim p^{v}$ for two dyadic numbers $p^{u},p^{v}\geq 1$. We can also assume $p^{u}p^{v}\gtrapprox\#E$. Hence, assuming Conjecture 4, we have
@@
However, we still need the condition $p^{u}\gtrsim p^{s_{1}}$. This is handled by the following trick. We will conduct an algorithm to find a sequence $E=E_{1}\supset E_{2}\supset\cdots\supset E_{m}$, and disjoint subsets $\Xi_{1},\Xi_{2}\cdots,\Xi_{m}$ of $\mathbf{0}\times\mathbb{F}_{p}$. To start with, let $E_{1}=E$. By pigeonholing, we can find $\Xi_{1}\subset\mathbf{0}\times\mathbb{F}_{p}$, such that #$\#\Xi_{1}\sim p^{u_{1}}$, #$\#E(\xi)\sim p^{v_{1}}$ for each $\xi\in\Xi$, and
$p^{u_{1}}p^{v_{1}}\gtrsim_{\varepsilon}\#E_{1}\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{1},t_{2};2,1)-\varepsilon}.$
Suppose we have defined $E_{i},\Xi_{i}$ for $11\leq i\leq j-$ 1. If $\sum_{i=1}^{j-1}p^{u_{i}}\ll p^{s_{1}}$, then we define
@@
Since each $\ell\in\mathcal{L}$ is not parallel to $\mathbb{F}_{p}\times\mathbf{0}$, we see that $E_{j}$ is still a $(s_{1},t_{2};2,1)$-set. By the same pigeonholing, we find $\Xi_{j}\subset(\mathbf{0}\times\mathbb{F}_{p})\setminus\bigcup_{1\leq i\leq j-1}\Xi_{i}$, so that #$\#\Xi_{j}\sim p^{u_{j}}$, #$\#E(\xi)\sim p^{v_{j}}$ for each $\xi\in\Xi_{j}$, and
@@
If $\sum_{i=1}^{j-1}p^{u_{i}}\gtrsim p^{s_{1}}$, then we stop and let $m=j-1$.
The algorithm will stop in finite steps. Denote $\Xi:=\bigcup_{1\leq i\leq m}\Xi_{i}$. Let $p^{u},p^{v}$ be dyadic numbers so that $p^{v}=\min_{1\leq i\leq m}p^{v_{i}}$, $p^{u}\sim\#\Xi$. For each $\xi\in\Xi$, let $Y_{\xi}$ be a subset of $E(\xi)$ with #$\#Y_{\xi}\sim p^{v}$. This is doable, since #$\#E(\xi)\gtrsim p^{v}$ if $\xi\in\Xi$. One can check
@@
Return back to $\widetilde{E}_{W}:=\bigcup_{V\in\mathcal{V}_{W}}Y(\ell_{V})$ (which lie in $\widetilde{\mathbb{F}_{p}^{2}}=\mathbb{F}_{p}\times\mathbf{0}^{k-1}\times\mathbb{F}_{p}$). Recall #$\#\mathcal{V}_{W}\gtrapprox p^{t_{2}}$ and #$\#Y(\ell_{V})\sim p^{s_{1}}$. We see that $\widetilde{E}_{W}$ is a $(s_{1},(t_{2}-\varepsilon)_{+};2,1)$-set, for any $\varepsilon>0$. Here, $(t_{2}-\varepsilon)_{+}=\max\{t_{2}-\varepsilon,0\}$.
By Lemma 31, there exist dyadic numbers $p^{u(W)},p^{v(W)}$, so that $\widetilde{E}_{W}$ contains a subset of form
@@
Here, $\Xi_{W}\subset\mathbf{0}^{k}\times\mathbb{F}_{p}$ with #$\#\Xi_{W}\sim p^{u(W)}$; $\rho_{\xi}:=\xi+(\mathbb{F}_{p}\times\mathbf{0}^{k})$ and $Y_{W}(\rho_{\xi})\subset\rho_{\xi}$ with #$\#Y_{W}(\rho_{\xi})\sim p^{v(W)}$. We also have
@@
By pigeonholing on $(u(W),v(W))$ again, we can throw away some $W\in\mathcal{W}$ so that the cardinality of $\mathcal{W}$ is still $\gtrapprox p^{t_{1}}$, and assume there exist $u,v$ such that
[…]
@@
(35) $\begin{split}\bigcup_{V\in\mathcal{V}_{W}}Y(V)&=\bigcup_{V\in\mathcal{V}_{W}}\bigsqcup_{x\in Y(\ell_{V})}Y(V,x)\\ &=\bigcup_{x\in\widetilde{E}_{W}}\bigg{(}\bigcup_{V:V\in\mathcal{V}_{W},x\in Y(\ell_{V})}Y(V,x)\bigg{)}.\end{split}$
@@
Then $Y(W,x)$ is contained in the $(k-1)$-plane $x+W$.
@@
The reader can check, for fixed $\xi\in\Xi_{W}$, $\bigcup_{x\in Y_{W}(\rho_{\xi})}Y(W,x)$ lies in $\mathbb{F}_{p}^{k}\times\{\xi\}$.
@@
We want to estimate the lower bound of the right hand side. Intuitively, one can imagine the worst case is when $\Xi_{W}$ are morally the same for all $W\in\mathcal{W}$; otherwise we may translate them to the same $(k+1)$-th coordinate so that their union becomes smaller. For example, if we have sets $A_{i}\times\{\xi_{i}\}\subset\mathbb{F}_{p}^{k}\times\mathbb{F}_{p}$ ($i\in I$). Then $\bigcup_{i\in I}A_{i}\times\{\xi_{i}\}$ is the smallest when all the $\xi_{i}$ are the same.
@@
#$\#E^{\prime}\geq\bigcup_{\xi\in\Xi_{W}}\bigcup_{x\in Y_{W}(\rho_{\xi})}Y(W,x)\gtrsim p^{u}p^{v}p^{s_{2}}.$
@@
(36) $p^{t_{\xi}}\leq\#\mathcal{W}\cdot\#Y_{W}(\rho_{\xi})\lesssim p^{t_{1}}p^{v};$
(37) $\sum_{\xi\in\mathbf{0}^{k}\times\mathbb{F}_{p}}p^{t_{\xi}}=\sum_{W\in\mathcal{W}}\#\Xi_{W}\cdot\#Y_{W}(\rho_{\xi})\gtrapprox p^{t_{1}}p^{u}p^{v}.$
We used that $p^{t_{1}}\gtrsim\#\mathcal{W}\gtrapprox p^{t_{1}}$, #$\#\Xi_{W}\sim p^{u}$, #$\#Y_{W}(\rho_{\xi})\sim p^{v}$.
And note that $Y(W,x)$ is a subset of the line $x+W$ such that #$\#Y(W,x)\gtrsim p^{s_{2}}$. Therefore, $E_{\xi}^{\prime}$ is a $(s_{2},t_{\xi};k,k-1)$-set. By (27), we have
#$\#E^{\prime}_{\xi}\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{2},t_{\xi};k,k-1)-\varepsilon}.$
@@
#$\#E\gtrsim_{\varepsilon}\sum_{\xi\in\mathbf{0}^{k}\times\mathbb{F}_{p}}p^{\mathbf{F}(s_{2},t_{\xi};k,k-1)-\varepsilon}.$
We claim that the right hand side is
@@
#$\#\Xi_{\circ}\cdot p^{t_{\circ}}\gtrapprox p^{t_{1}}p^{u}p^{v}.$
@@
#$\#\Xi_{\circ}\gtrapprox p^{u}p^{t_{1}+v-t_{\circ}}.$

### v2

@@ 4 Furstenberg set problem: Lower bound I
Suppose $E=\bigcup_{V\in\mathcal{V}}Y(V)$. Note that each $x\in E$ can lie in at most $G(k,\mathbb{F}_{p}^{n})\sim p^{k(n-k)}$ different $V$. We have
$E\cdot p^{k(n-k)}\geq\sum_{V\in\mathcal{V}}\#Y(V)\gtrsim p^{t}.$
Hence, $E\gtrsim p^{t-k(n-k)}$. Also, trivially, $E\geq 1$. ∎
Proposition 4.2.
@@
By the definition of $(s,0;n,k)$-set, $E\geq\#Y(V)\gtrsim p^{s}$. ∎
In this section, we prove Theorem 1.8 for $(n,k)=(k+1,k)$. The proof is by induction on $k$. The goal of this section is to prove the following result.
Proposition 4.3.
Let $k\in\mathbb{N}$ and $k\geq 2$. Suppose for admissible (s,t;k,k1)(s,t;k,k-1), we know
$\inf_{E:\ (s,t;k,k-1)\textup{-set}}\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;k,k-1)-\varepsilon}$ (27)
@@
$\inf_{E:\ (s,t;k+1,k)\textup{-set}}\#E\gtrsim_{\varepsilon}p^{\mathbf{F}(s,t;k+1,k)-\varepsilon},$ (28)
@@
We mention a technical thing in the proof. For $\eta>0$, denote $\eta\mathbb{Z}:=\{\eta n:n\in\mathbb{Z}\}$. In Theorem 1.8, the implicit constant in the inequality also depends on $s,t$, and therefore may not be uniformly bounded. To get rid of it, we choose a small $\eta>0$, and discretize $s,t$, so that $s,t\in\eta\mathbb{Z}$. Then we obtain the bound
$E\gtrsim_{\varepsilon,\eta,n,k}p^{\mathbf{F}(s-\eta,t-\eta;n,k)-\varepsilon},$
where the implicit constant is independent of $s,t$.
Lemma 4.4.
Assume (27) holds. Suppose $(s,t;k+1,k)$ is admissible. Fix $0<\varepsilon<1/100$. Then for any $(s,t;k+1,k)$-set $E$, there exist numbers $t_{1},t_{2},s_{1},s_{2},u,v$, such that
$\begin{cases}1\leq p^{t_{1}}\leq p^{k-1},1\leq p^{t_{2}}\leq p^{2};\\ 1\leq p^{s_{1}}\leq p,1\leq p^{s_{2}}\leq p^{s};\\ p^{s_{1}}\lesssim p^{u}\leq p,1\leq p^{v}\leq p;\end{cases}$ (29)
@@
$\begin{cases}p^{t_{1}+t_{2}}\gtrapprox p^{t};\\ p^{s_{1}+s_{2}}\gtrapprox p^{s};\\ p^{u+v}\gtrsim_{\varepsilon}p^{\mathbf{F}((s_{1}-\varepsilon^{2})_{+},(t_{2}-\varepsilon^{2})_{+};2,1)-\varepsilon};\end{cases}$ (30)
@@
$E\gtrsim_{\varepsilon}p^{u+\max\{\mathbf{F}((s_{2}-\varepsilon^{2})_{+},(t_{1}+v-\varepsilon^{2})_{+};k,k-1),s_{2}+v\}-\varepsilon}.$ (31)
Here, $(x)_{+}=\max\{0,x\}$. In addition, the implicit constants do not depend on $t_{1},t_{2},s_{1},s_{2},u,v$.
@@
We will do several steps of pigeonholings, which lead to the parameters $t_{1},t_{2},s_{1},s_{2},u,v$.
@@
where $\mathcal{V}\subset A(k,\mathbb{F}_{p}^{k+1})$ with $\mathcal{V}\sim p^{t}$, and $Y(V)\subset V$ with $Y(V)\sim p^{s}$.
@@
Now, we partition $\mathcal{V}$ according to $G(k-1,\mathbb{F}_{p}^{k}\times\mathbf{0})$. Since each $V\in\mathcal{V}$ is not parallel to $\mathbb{F}_{p}^{k}$ and the ambient space is $\mathbb{F}_{p}^{k+1}$, $V$ must intersect $\mathbb{F}_{p}^{k}\times\mathbf{0}$ at a (k1)(k-1)-plane. For each $W\in G(k-1,\mathbb{F}_{p}^{k}\times\mathbf{0})$, we define $\mathcal{V}_{W}$ to be the set of $V\in\mathcal{V}$ such that $V\cap\mathbb{F}_{p}^{k}$ is parallel to $W$. We obtain a partition
@@
First pigeonholing. We assume that there exists $\mathcal{W}\subset G(k-1,\mathbb{F}_{p}^{k}\times\mathbf{0})$ with $\mathcal{W}\sim p^{t_{1}}$ such that for each $W\in\mathcal{W}$, $\mathcal{V}_{W}\sim p^{t_{2}}$. Here, $1\leq p^{t_{1}}\leq p^{k-1},1\leq p^{t_{2}}\leq p^{2}$ are dyadic numbers, and
$p^{t_{1}}p^{t_{2}}\gtrapprox p^{t}.$ (32)
(Recall that $A\gtrapprox B$ means $A\gtrsim(\log p)^{C}B$.) We still use $\mathcal{V}$ to denote $\bigcup_{W\in\mathcal{W}}\mathcal{V}_{W}$. And we see that
$\mathcal{V}\gtrapprox p^{t}.$
Second pigeonholing. By properly choosing the coordinate and throwing away $\lesssim 1$ fraction of $W$ in $\mathcal{W}$, we can assume that each $W$ in $\mathcal{W}$ satisfies $W\cap(\mathbb{F}_{p}\times\mathbf{0}^{k})=\mathbf{0}^{k+1}$. Denote $\widetilde{\mathbb{F}_{p}^{2}}=\mathbb{F}_{p}\times\mathbf{0}^{k-1}\times\mathbb{F}_{p}$. If $V\in\mathcal{V}_{W}$ and $W\in\mathcal{W}$, then $V\cap\widetilde{\mathbb{F}_{p}^{2}}$ is a line. We denote it by $\ell_{V}:=V\cap\widetilde{\mathbb{F}_{p}^{2}}$. Note that $\ell_{V}$ is not parallel to the first coordinate in $\widetilde{\mathbb{F}_{p}^{2}}$.
@@
Noting that $Y(V)\sim p^{s}$, by pigeonholing, we find a subset $Y(\ell_{V})\subset\ell_{V}$, so that $Y(\ell_{V})\sim s_{1}(V)$, and for each $x\in Y(\ell_{V})$, $Y(V)\cap(x+W)\sim s_{2}(V)$. Here, $1\leq p^{s_{1}(V)}\leq p,1\leq p^{s_{2}(V)}\leq p^{s}$ are dyadic numbers that satisfy
@@
$Y(V,x)$ is called the horizontal $x$-slice. (Note that there is no confusion since $V$ uniquely determines $W$.) Thus, we have
$Y(V)=\bigsqcup_{x\in Y(\ell_{V})}Y(V,x).$ (33)
@@
We summarize what we have done so far. We can redefine $\mathcal{W}$, $\mathcal{V}_{W}$ and $\mathcal{V}$, so that $\mathcal{V}=\bigcup_{W\in\mathcal{W}}\mathcal{V}_{W}$ and the following uniformity hold.
$\mathcal{W}\gtrapprox p^{t_{1}}.$
$\mathcal{V}_{W}\gtrapprox p^{t_{2}},\textup{~for~}W\in\mathcal{W}.$
@@
$Y(V)=\bigsqcup_{x\in Y(\ell_{V})}Y(V,x),$ (34)
with $Y(\ell_{V})\sim p^{s_{1}}$ and for each $x\in Y(\ell_{V})$, $Y(V,x)\sim p^{s_{2}}$.
[…]
Lemma 4.5.
Assume Conjecture 1.5 holds. Suppose $E=\bigcup_{\ell\in\mathcal{L}}Y(\ell)$ is a $(s_{1},t_{2};2,1)$-set. In addition, we assume that each $\ell\in\mathcal{L}$ is not parallel to $\mathbb{F}_{p}\times\mathbf{0}$. Then there exist dyadic numbers $p^{u},p^{v}$ such that
@@
for any $\varepsilon>0$. In addition, $E$ contains a subset of the form
@@
where $\Xi\subset\mathbf{0}\times\mathbb{F}_{p}$ with $\Xi\sim p^{u}$, and $Y_{\xi}\subset E\cap\bigg(\xi+(\mathbb{F}_{p}\times\mathbf{0})\bigg)$ with $Y_{\xi}\sim p^{v}$.
@@
For $\xi\in\mathbf{0}\times\mathbb{F}_{p}$, define $E(\xi):=E\cap\bigg(\xi+(\mathbb{F}_{p}\times\mathbf{0})\bigg)$. By pigeonholing on horizontal slices of $E$, we can find $\Xi\subset\mathbf{0}\times\mathbb{F}_{p}$ so that for all $\xi\in\Xi$, $E(\xi)$ have comparable cardinality. We may denote $\Xi\sim p^{u},\#E(\xi)\sim p^{v}$ for two dyadic numbers $p^{u},p^{v}\geq 1$. We can also assume $p^{u}p^{v}\gtrapprox\#E$. Hence, assuming Conjecture 1.5, we have
@@
However, we still need the condition $p^{u}\gtrsim p^{s_{1}}$. This is handled by the following trick. We will perform an algorithm to find a sequence $E=E_{1}\supset E_{2}\supset\cdots\supset E_{m}$, and disjoint subsets $\Xi_{1},\Xi_{2}\cdots,\Xi_{m}$ of $\mathbf{0}\times\mathbb{F}_{p}$. To start with, let $E_{1}=E$. By pigeonholing, we can find $\Xi_{1}\subset\mathbf{0}\times\mathbb{F}_{p}$, such that $\Xi_{1}\sim p^{u_{1}}$, $E(\xi)\sim p^{v_{1}}$ for each $\xi\in\Xi$, and
$p^{u_{1}}p^{v_{1}}\gtrapprox\#E_{1}\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{1},t_{2};2,1)-\varepsilon}.$
Suppose that we have defined $E_{i},\Xi_{i}$ for $1\leq i\leq j-1$. If $\sum_{i=1}^{j-1}p^{u_{i}}\ll p^{s_{1}}$, then we define
@@
Since each $\ell\in\mathcal{L}$ is not parallel to $\mathbb{F}_{p}\times\mathbf{0}$, we see that $E_{j}$ is still a $(s_{1},t_{2};2,1)$-set. By the same pigeonholing, we find $\Xi_{j}\subset(\mathbf{0}\times\mathbb{F}_{p})\setminus\bigcup_{1\leq i\leq j-1}\Xi_{i}$, so that $\Xi_{j}\sim p^{u_{j}}$, $E(\xi)\sim p^{v_{j}}$ for each $\xi\in\Xi_{j}$, and
@@
If $\sum_{i=1}^{j-1}p^{u_{i}}\gtrsim p^{s_{1}}$, then we stop and let m=j1m=j-1.
The algorithm will stop in finite steps. Denote $\Xi:=\bigcup_{1\leq i\leq m}\Xi_{i}$. Let $p^{u},p^{v}$ be dyadic numbers so that $p^{v}=\min_{1\leq i\leq m}p^{v_{i}}$, $p^{u}\sim\#\Xi$. For each $\xi\in\Xi$, let $Y_{\xi}$ be a subset of $E(\xi)$ with $Y_{\xi}\sim p^{v}$. This is doable, since $E(\xi)\gtrsim p^{v}$ if $\xi\in\Xi$. One can check
@@
Return to $\widetilde{E}_{W}:=\bigcup_{V\in\mathcal{V}_{W}}Y(\ell_{V})$ (which lies in $\widetilde{\mathbb{F}_{p}^{2}}=\mathbb{F}_{p}\times\mathbf{0}^{k-1}\times\mathbb{F}_{p}$). Recall $\mathcal{V}_{W}\gtrapprox p^{t_{2}}$ and $Y(\ell_{V})\sim p^{s_{1}}$. We see that $\widetilde{E}_{W}$ is a $(s_{1},(t_{2}-\varepsilon)_{+};2,1)$-set, for any $\varepsilon>0$.
By Lemma 4.5, there exist dyadic numbers $p^{u(W)},p^{v(W)}$, so that $\widetilde{E}_{W}$ contains a subset of form
@@
Here, $\Xi_{W}\subset\mathbf{0}^{k}\times\mathbb{F}_{p}$ with $\Xi_{W}\sim p^{u(W)}$; $\rho_{\xi}:=\xi+(\mathbb{F}_{p}\times\mathbf{0}^{k})$ and $Y_{W}(\rho_{\xi})\subset\rho_{\xi}$ with $Y_{W}(\rho_{\xi})\sim p^{v(W)}$. We also have
@@
By pigeonholing on $(u(W),v(W))$ again, we can throw away some $W\in\mathcal{W}$ so that the cardinality of $\mathcal{W}$ is still $\gtrapprox p^{t_{1}}$, and assume that there exist $u,v$ such that
$(u(W),v(W))=(u,v),\ \textup{~for~all~}W\in\mathcal{W},$
@@
$\begin{split}\bigcup_{V\in\mathcal{V}_{W}}Y(V)&=\bigcup_{V\in\mathcal{V}_{W}}\bigsqcup_{x\in Y(\ell_{V})}Y(V,x)\\ &=\bigcup_{x\in\widetilde{E}_{W}}\bigg(\bigcup_{V:V\in\mathcal{V}_{W},x\in Y(\ell_{V})}Y(V,x)\bigg).\end{split}$ (35)
@@
Then $Y(W,x)$ is contained in the (k1)(k-1)-plane $x+W$.
@@
The reader can check that, for fixed $\xi\in\Xi_{W}$, $\bigcup_{x\in Y_{W}(\rho_{\xi})}Y(W,x)$ lies in $\mathbb{F}_{p}^{k}\times\{\xi\}$.
@@
We want to estimate the lower bound of the right-hand side. Intuitively, one can imagine that the worst case is when $\Xi_{W}$ are morally the same for all $W\in\mathcal{W}$; otherwise we may translate them to the same $(k+1)$-th coordinate so that their union becomes smaller. For example, if we have sets $A_{i}\times\{\xi_{i}\}\subset\mathbb{F}_{p}^{k}\times\mathbb{F}_{p}$ ($i\in I$). Then $\bigcup_{i\in I}A_{i}\times\{\xi_{i}\}$ is the smallest when all the $\xi_{i}$ are the same.
@@
$E^{\prime}\geq\bigcup_{\xi\in\Xi_{W}}\bigcup_{x\in Y_{W}(\rho_{\xi})}Y(W,x)\gtrsim p^{u}p^{v}p^{s_{2}}.$
@@
$p^{t_{\xi}}\leq\#\mathcal{W}\cdot\#Y_{W}(\rho_{\xi})\lesssim p^{t_{1}}p^{v};$ (36)
$\sum_{\xi\in\mathbf{0}^{k}\times\mathbb{F}_{p}}p^{t_{\xi}}=\sum_{W\in\mathcal{W}}\#\Xi_{W}\cdot\#Y_{W}(\rho_{\xi})\gtrapprox p^{t_{1}}p^{u}p^{v}.$ (37)
We used that $p^{t_{1}}\gtrsim\#\mathcal{W}\gtrapprox p^{t_{1}}$, $\Xi_{W}\sim p^{u}$, $Y_{W}(\rho_{\xi})\sim p^{v}$.
And note that $Y(W,x)$ is a subset of the line $x+W$ such that $Y(W,x)\gtrsim p^{s_{2}}$. Therefore, $E_{\xi}^{\prime}$ is a $(s_{2},t_{\xi};k,k-1)$-set. By (27), we have
$E^{\prime}_{\xi}\gtrsim_{\varepsilon}p^{\mathbf{F}(s_{2},t_{\xi};k,k-1)-\varepsilon}.$
@@
$E\gtrsim_{\varepsilon}\sum_{\xi\in\mathbf{0}^{k}\times\mathbb{F}_{p}}p^{\mathbf{F}(s_{2},t_{\xi};k,k-1)-\varepsilon}.$
We claim that the right-hand side is
@@
$\Xi_{\circ}\cdot p^{t_{\circ}}\gtrapprox p^{t_{1}}p^{u}p^{v}.$
@@
$\Xi_{\circ}\gtrapprox p^{u}p^{t_{1}+v-t_{\circ}}.$